# Recipes for Pre-training LLMs with MXFP8

### v1

@@ Abstract
Precision scaling — using fewer bits to represent model parameters and related tensors during pre-training — has emerged as a compelling technique for improving GPU efficiency without sacrificing accuracy. Microscaling (MX) formats [1] in NVIDIA’s latest Blackwell GPUs represent a major leap in enabling this precision scaling aspect. These formats combine narrow floating-point data types with per-block scaling factors, offering a fine-grained approach to quantizing tensors.
Although MX-formats offer the promise of improved numeric stability compared to other reduced-precision representations, in practice they must be used carefully in order to successfully converge an LLM on a multi-trillion token dataset. In this paper, we show that the rounding mode suggested in OCP [1] specification can lead to divergence when pre-training an LLM. We show an improved rounding mode, which uses round-to-infinity to compute scaling factors, enables successful pre-training in MXFP8 for an 8B model on 15T tokens.
@@ 1 Introduction
Scaling the number of parameters of deep learning models, especially large generative models, has intensified the need for more efficient compute and memory solutions. Precision scaling — reducing number of bits to represent a data type — is a successful strategy to improve performance while lowering memory requirements. However, precision scaling without sacrificing model accuracy remains a key challenge.
Microscaling (MX) formats, developed in the Open Compute Project (OCP) [1], aims to offer a standardized approach to low-precision representation. By combining narrow floating-point types with fine-grained scaling factors, MX-formats aim to deliver a balance of dynamic range, precision, and hardware efficiency. OCP [1] and research works like [2, 3, 4] have shown a fine-grained scaling approach to be an effective strategy compared to granular (e.g., per-tensor) scaling approaches, especially in sub-8-bit pre-training regimes.
NVIDIA’s latest GPU generation, Blackwell [5], adds native MX data type support to Tensor Cores. Blackwell supports MX data for 8-bit (MXFP8), 6-bit (MXFP6) and 4-bit (MXFP4) floating-point types. This paper primarily focuses on MXFP8 and provides an in-depth analysis of pre-training large language models (LLMs). It details the conversion process from high precision formats to the quantized MXFP8 format, including rounding modes for data and scaling factors. These insights aim to guide the efficient and accurate pre-training of LLMs using MXFP8.
Notably, we observe that the rounding mode suggested in OCP [1] for scale factor computation causes MXFP8 pre-training to differ from a BF16 baseline. We propose a modification to this scale factor rounding where we round-to-infinity instead of rounding-down, and demonstrate that this adjustment enables MXFP8 pre-training to match BF16 pre-training accuracy.
The primary contributions of this paper are as follows:
  • A methodical study of MX quantization for LLM pre-training where we investigate the following design choices: data formats, quantization schemes, storage, the selection of tensors to quantize, and examining how these choices affect accuracy. Specifically, we identify E4M3 to be the MXFP8 data type that best helps maintain pre-training accuracy.
  • We find that the rounding mode suggested in the OCP specification can lead to divergence during pre-training. Building on the analysis, we propose a rounding mode for scale factors that lets MXFP8 trained models to match BF16 model accuracy. Specifically, we describe a recipe to pretrain LLMs with MXFP8 and show experiments where an 8 billion parameter dense LLM trained on 15 trillion tokens matches a BF16 trained model and a per-tensor scaled FP8 [6] model with first and last layer in BF16.
For each design choice, we present ablation studies on small models (e.g. 100s of millions of parameters) pretrained on short (e.g. 100s of billions) token horizons. We then apply the findings to large multi-billion parameter models trained on trillion token horizons. This makes our conclusions robust for large-scale foundation model pre-training.
@@ 2 Microscaling format support in NVIDIA Blackwell
Background: An MX-format is specified by a block size K, a shared scaling factor per block, X, and the data-type of elements in the block. A block is a contiguous set of K elements along an axis in a tensor. K = 32 for all MX types on Blackwell. The data-type of X is UE8M0, an unsigned exponent-only value that encodes either NaN or any power-of-two [^1] Power-of-two is a number of the form $2^{n}$ where n is a positive integer, negative integer, or zero. value in range $2^{-127}$ to $2^{127}$.
Given K input elements in a source-format, $V_{i}$ (typically FP32); $1\leq i\leq K$, conversion to MX-format consists of computing X and $Q_{i}$ such that $Q_{i}\times$X is the decoded value corresponding to $V_{i}$. X and $Q_{i}$ are stored in memory instead of $V_{i}$. This is depicted in Fig. 1. A tensor in source-format is sub-divided into blocks of K elements and converted into MX-format to be stored in memory and/or processed by math units in hardware. Tensor Cores in Blackwell consume X and $Q_{i}$ to compute the dot product of two MX-formatted blocks. If the accumulated output of a dot-product operation in Tensor Cores is FP32, then this output is thereafter quantized to MX-format if a subsequent operation consuming the output needs it in that format. The conversion process is, $Q_{i}=\verb+Quantize_to_fp8+(V_{i}/2^{X})$. Section 3 describes the conversion details. Fine-grained scaling helps each MX block independently align to the needed range of input values before quantization.
Table 1 shows the MX-formats supported in Blackwell. The data type column uses the convention ExMy to denote x-bit for floating-point exponent and y-bit for mantissa. For a fixed bit-width, floating-point numbers trade-off exponent width with mantissa width. Thus, MXFP8 E4M3 is a 8-bit data type with with 1 sign bit, 4-bit for FP exponent and 3-bit of mantissa. Similarly, E5M2 is an 8-bit type with 5-bit for exponent and 2-bit for mantissa. Compared to E4M3, E5M2 can represent a larger dynamic range, but with less precision. E5M2 is the only type that follows IEEE 754 conventions [7] for representation of special values. For all other data types, dynamic range is extended by not representing Infinity and NaN (E4M3 has only one bit-pattern for NaN [6]). More bits in the exponent field translates to a larger range while more bits in the mantissa field translates to more precision within a given range. Every floating-point type has a dynamic range that it can represent — we denote this range in terms of binades which is the log2-ratio of the maximum to the minimum finite representable in that format.
@@ 3 Pre-training with MXFP8
Model weights, activations, and gradient tensors are quantized from FP32 into MXFP8 during the forward and backward passes in training. MXFP8 tensors are then stored and operated on by the hardware. We first describe the conversion process, Quantize_to_fp8$(V_{i}/2^{X})$, starting with scale computation and following on to the quantization procedure for each element. The quantization process described below is the same across all MX-formats (E4M3, E5M2 , E2M3, E3M2, and E2M1), with only the MX data type differing.
3.1 Conversion from FP32 to MXFP8
Computing X: Typically, most of the values within each block in a tensor will be out of the representable range for the target MX format, both underflowing the minimum and overflowing the maximum representable number. To address this, a scale factor is multiplied with all values within a block, to shift the maximum number of values to be within the representable range.
The scale factor, X, is computed based on the absolute maximum value (amax) among the 32 high-precision input values, i.e. amax = $\max\|V_{i}\|;1\leq i\leq 32$. The goal is to map this amax in the input-format to be the largest representable value in the desired MX-format. Special care is taken if some or all the elements in the input are Infinity and/or not-a-number (NaN). If amax in a block is 0, then X is set to -127, thus, interpreted value of X is $2^{-127}$. All $Q_{i}$’s in this case are set to 0.
When X is not Inf, NaN or 0, the OCP [1] specification suggests X be set to the largest power-of-two less than or equal to amax divided by the largest power-of-two representable in the MX-format type. For example, with the E4M3 type, $X+=\frac{\text{floor}(\log_{2}(\texttt{amax}))}{\text{floor}(\log_{2}(448))}$, since 448 is the largest magnitude in E4M3. Thus, the OCP specification ignores the floating-point significand of this ratio. [^2] Assume amax and max representable in MX-format (destamax) are normal floating-point numbers of the form amax = $2^{A}$ 1.$M_{a}$ and destmax = $2^{E}$ 1.$M_{e}$ (where mantissa $M_{a}$ and $M_{e}$ lie between 0 (inclusive) and 1. OCP rounding scheme results in scale exponent = A-E (ignoring the case where min(A-E) is clamped to -127)
We observe accuracy degradation when following the OCP specification. Fig. 2 show training loss curves for a 843 million parameter transformer model trained under two token horizons: 300 billion and 1 trillion tokens. Two different configurations are studied:
  • (cfg1) E4M3 for all tensors (weights (W), activations (A) and gradients (G)), and
  • (cfg2) E4M3 for W and A tensors, E5M2 for G tensors.
The E5M2 format has $\sim$1.8x more binades than E4M3 and since gradients typically have larger dynamic range, [6] advocated the use of E5M2 with per-tensor scaling method. This rationale underpins the study of E5M2 gradients. In both cfg1 and cfg2, scale factor computation using OCP method leads to training divergence (Fig. 2(a)) or a widening loss gap relative to BF16 (Fig. 2(b)).
Figure 2: Comparing scale factor rounding modes suggested in the OCP specification and our work. E4M3 and E5M2 are MXFP8 formats; notation E4M3(W) denotes E4M3 data type for Weights. Similarly, E4M3(A) and E4M3 (G) refer to activations and gradients, respectively.
Algorithm 1 outlines our method for computing the scale factor. The key change is to round-up towards positive infinity the exponent of ratio of amax and max representable in MX-format, destmax (while saturating to UE8M0 max/min limits). This is in contrast to the OCP scheme which effectively suggests to round-down the scale value. Since a high-precision value, $V_{i}$, is scaled by the scale factor $2^{X}$, rounding-up the denominator of the fraction ($V_{i}/2^{X}$) has a tendency to map amax below destmax. In contrast, the OCP method has a tendency to map amax above destmax (which subsequently needs to be clamped so that it becomes representable). We hypothesize that this saturating effect with OCP rounding method affects model accuracy. A more detailed and accurate workflow to compute and round the value of X is described in Appendix Sec. A.1.
$\text{expX}_{\text{float}}\leftarrow\log_{2}(X_{\text{float}})$triangleright extract the exponent of X (de-biased form)
$\text{expX}_{\text{int}}\leftarrow\text{ceil}(\text{expX}_{\text{float}})$triangleright round-up
$X\leftarrow\text{clamp}(\text{expX}_{\text{int}},-127,127)$triangleright clamp to min/max E8M0 representable
$X\leftarrow X+127$triangleright add bias
return $X$triangleright store X
Fig. 2 shows that MXFP8 with E4M3 for gradients (blue curve) and MXFP8 with E5M2 for gradients (purple curve) with our proposed rounding scheme both overlap with a reference BF16 loss curve across both 300B and 1T token horizons. Later, in section 3.2 we discuss that E4M3 is in fact a better choice for pre-training LLMs versus E5M2 with MX style fine-grained scaling.
Quantizing FP32 values to MX type: Once X is computed, $V_{i}$ is scaled by $2^{X}$ and the resulting value is quantized to a FP8-representable number. This is the Quantize_to_fp8() function. Round-to-nearest-ties-to-even (RN) rounding is used during this quantization step. The conversion process is saturating, i.e. if after rounding the resulting value exceeds FP8 max or is less than FP8 min value, then the result is clamped to respective max or min value.
A practical instance of this conversion process in low-precision LLM pre-training arises when matrix multiplication and accumulation (MMA) outputs, typically stored in FP32, must be mapped to MXFP8. In this case 8-bit quantized values are stored in memory, thus saving write bandwidth and storage capacity when compared with storing FP32 values. Subsequent model operation then reads MXFP8 values, saving read bandwidth compared to loading FP32 values. Further, since Tensor Cores can process MX-formatted inputs, the MMA operation in lower precision consumes less energy and operates at higher throughput.
3.2 E4M3 data type for weights, activations and gradients
@@
FP8 has two variants in Blackwell 1: E4M3 and E5M2. Our experiments show the following:
  • Better convergence with E4M3 instead of E5M2 for weights and activations during training. Fig. 3(a) shows the loss behavior comparing E4M3 and E5M2 for the same 843M model used in Fig. 2(b). The purple curve has E5M2 for activations and the blue curve has E5M2 for weights while all other tensors are in E4M3. Both these curves show worse loss convergence compared to using E4M3 for all tensors (orange curve) or using E5M2 for gradients (yellow curve).
  • E4M3 for gradients maintains training loss parity with BF16 pre-training, especially for models with 2 billion or more parameters. Fig. 3(b) shows the loss behavior comparing E4M3 and E5M2 for the gradient tensor on an 8 billion parameter LLM trained on 1 trillion tokens: using E4M3 gradients (orange) has lower loss than using E5M2 gradients (yellow). This gap increases as the model is trained on more tokens. This change in behavior with increasing model parameter counts underscores the importance of examining the numerical properties of formats across a broad range of model sizes.
Previously, per-tensor scaled FP8 studies in [8, 6, 9] and coarse-grained block-scaled FP8 studies in [10] used E5M2 for gradient tensors in lieu of E4M3. With fine-grained scaling, the dynamic range requirements at a 32-element block size is sufficiently captured by 17.8 binades in E4M3 type. Once the range requirements are met, precision (or sampling) becomes important, and E4M3 having 8 samples per binade is better than E5M2 having only 4. Thus, for our MXFP8 pre-training recipe, we quantize all three tensor types – weights, activations, and gradients – with the E4M3 data type.
Figure 4: Top: Transformer layers quantized to MXFP8 inside a single transformer block. Bottom: Training workflow for a single layer during FPROP, DGRAD and WGRAD.
Layers to instantiate in MXFP8 and training workflow: We use language-based transformer models for all studies in this paper. Our future work will study these recipes on speech and vision models. Based on our studies, our guideline is to quantize QKV, PROJ and FFN Up- and Down-projection layers to MXFP8 across all transformer blocks in the model. The Batch-Matrix Multiplications (BMM1, the query-key dot product and BMM2, the attention score-value product) in the self-attention layer, along with operations like Softmax, Act-func and residual-add are in high-precision. We found this to be the safest option for maintaining accuracy parity with BF16 pre-training. This is illustrated in Fig. 4. The input embedding layer and the final output-projection layer are also in BF16 or FP16. All studies in this paper use this guideline.
During training, with MXFP quantization, the training framework has to keep two copies of each of the tensors – weights, activations and gradients: each copy is quantized along the axes of dot-product reduction (row and column). Fig. 4 shows how each tensor is used in forward (FPROP), weight-gradient (WGRAD) and activation-gradient (DGRAD) computation during the training loop. Since each tensor is used in non-transposed and transposed form, quantization needs to occur along two separate axes (row and column).
Summary so far: We introduce a new rounding scheme for the MX scale factor that addresses the divergence caused by the OCP-based approach and achieves loss parity with BF16 on an 843M-parameter model up to 1T tokens. Additionally, adopting E4M3 and our scale factor calculation method in Alg. 1 enables scaling to an 8B-parameter model trained on 15T tokens — representing the largest LLM pre-training at this scale with MXFP formats, to the best of our knowledge. Appendix (section A.4) shows that our recipe also holds for a 16B-parameter mixture-of-experts model.
3.3 Results with MXFP8 pre-training a LLM on 15T tokens
Figure 5: Pre-training a 8B LLM on 15T tokens. Top: Training behavior of BF16 vs MXFP8. Bottom: Comparing BF16, FP8 and MXFP8’s downstream task scores on MMLU and a set of 9 reasoning tasks. MXFP8 numerics uses our proposed rounding method and E4M3 for all quantized tensors.
We pre-train an 8B parameter Nemotron model [11] using Megatron-LM [12]. The model has 32 transformer blocks, 32 attention heads, hidden size is 4096, GQA group size is 8, KV-channels is 128, sequence length during pre-training is 8192. It is trained on 15T tokens with a batch size of 768. Initial learning rate is 6e-4 that cosine decays to 6e-6. A phased data-blending approach is used to train the model: in the first phase, a data mixture that promotes diversity in data is used and in the second phase high-quality datasets (e.g., Wikipedia) are used. We switch to the second phase at the 60% point of training. This blend style has also been used in other large scale pre-training setups [8]. The model is pre-trained on 3072 Hopper GPUs (since hardware with MX support was unavailable during much of the experimentation duration).
We use an emulation of MX-formats on Hopper GPUs: tensors that feed a MMA operation are first quantized into the MX-format and then cast back to BF16 before the BF16 MMA operation. The training workflow depicted in Fig. 4 is implemented in Megatron-LM. We validated the numerical fidelity of our emulation package by comparing with a 2B parameter LLM pre-training run on Blackwell using actual MXFP8 format and confirmed that they were identical.
Fig. 5 shows training loss behavior and task-level accuracy for the 8B pre-trained model. We report evaluation scores on two sets of downstream tasks: (1) 5-shot score on MMLU [13] and (2) Averaged 1-shot score across 9 general reasoning benchmarks: ARC-Challenge and ARC-Easy [14], Race [15], PIQA [16], Winogrande [17], Hellaswag [18], OpenBookQA [19], Social IQA [20] and Commonsense QA [21]. We observe the following:
  1. Validation perplexity of the model when pre-trained with MXFP8 matches pre-training with BF16 (left plot in Fig. 5). There is less than 0.50% difference between MXFP8 and BF16 validation perplexity values throughout the pre-training run.
  2. The middle and right plots in Fig. 5 show evaluation scores on two sets of downstream tasks. Again, scores for the MXFP8 trained model match with those of the BF16 trained model. This makes MXFP8 a viable candidate for pre-training LLMs.
MXFP8 versus FP8: In addition to MXFP8 and BF16, Fig. 5 also shows task-level scores for the same model trained with traditional FP8 precision. FP8 recipe uses software-emulated block scaling [6] where an entire tensor is scaled so that maximum number of tensor values fall into the representable range in the quantized format. We follow the guidelines suggested in [8] for FP8 pre-training setup: the first and last transformer block in the model are kept in BF16 while linear layers of the remaining blocks are quantized to FP8.  [8] found this choice appropriate for pre-training 8B and 56B parameter LLMs on 20T tokens. Keeping some layers in BF16 affects end-to-end speedup and also complicates pre-training — a choice has to be made on which layers to leave in higher-precision. We observe that MXFP8 matches FP8 accuracy on these two sets of tasks without requiring any BF16 layers.
MXFP8 versus blockwise-FP8: Further, some works like Deepseek-V3 [10] report the need to scale to a smaller block size when using FP8. In this setup, certain tensors require vector 1x128 scaling and certain tensors require per-block (e.g. 128x128) software scaling. This complicates GEMM kernel design. Native support for MXFP8 simplifies this — fine-grained scaling provides better numerical robustness and avoids any tradeoff between smaller block sizes and hardware speed.
In summary, we find MXFP8 maintains accuracy compared to BF16 or FP8 pre-trained models. On GB200 Blackwell systems, MXFP8 has 2$\times$ higher throughput than BF16 making end-to-end MXFP8 pre-training faster than BF16 pre-training. We also find the MXFP8 recipe to be simpler to use when compared to FP8 (all layers can be quantized and scaling is handled in the hardware) while allowing for equal or better throughput.

### v2

@@ Abstract
Using fewer bits to represent model parameters and related tensors during pre-training has become a required technique for improving GPU efficiency without sacrificing accuracy. Microscaling (MX) formats [1] introduced in NVIDIA Blackwell generation of GPUs [2] represent a major advancement of this technique, making it practical to combine narrow floating-point data types with finer granularity per-block scaling factors. In turn, this enables both quantization of more tensors than previous approaches and more efficient execution of operations on those tensors.
Effective use of MX-formats requires careful choices of various parameters. In this paper we review these choices and show how MXFP8-E4M3 datatype and a specific number conversion algorithm result in training sessions that match those carried out in BF16. We present results using models with up to 8B parameters, trained on high-quality datasets of up to 15T tokens.
@@ 1 Introduction
Scaling the number of parameters of deep learning models, especially large generative models, has intensified the need for more efficient compute and memory solutions. Reducing the number of bits to represent a data type (sometimes referred to as reducing precision) is a successful strategy to improve performance while lowering memory requirements. However, maintaining model accuracy while reducing precision remains a key challenge.
Microscaling (MX) formats, specified by an Open Compute Project working group (OCP) [1], combine narrow floating-point types with fine-grained scaling factors to deliver a balance of dynamic range, precision, and hardware efficiency. OCP v1.0 specification [1] and research works like [3, 4, 5] have shown a fine-grained scaling approach to be an effective strategy compared to granular (e.g., per-tensor) scaling approaches, especially in sub-8-bit pre-training regimes.
NVIDIA’s latest GPU generation, Blackwell [2], adds native MX data type support to Tensor Cores. Blackwell supports MX data for 8-bit (MXFP8), 6-bit (MXFP6) and 4-bit (MXFP4) floating-point types. This paper focuses on MXFP8 pre-training of large language models (LLMs), demonstrating two key factors needed for successful pre-training (no accuracy degradation compared to BF16 sessions):
  • First, MXFP8-E4M3 datatype should be used for all tensor types, including activation gradients.
  • Second, careful consideration of rounding is needed when converting from high-precision formats to MXFP8, resulting in an algorithm different from the one suggested in OCP v1.0 specification.
For each design choice, we present ablation studies on small models (e.g. 100s of millions of parameters) pretrained on short (e.g. 100s of billions) token horizons. We then apply the findings to large multi-billion-parameter models trained on trillion token horizons.
@@ 2 Microscaling format support in NVIDIA Blackwell
Background: An MX-format is specified by the block size K, a shared scaling factor per block, X, and the data-type of elements in the block. A block is a contiguous sequence of K elements. K = 32 for all MX types. The data-type of X is UE8M0 and X encodes either NaN or any power-of-two [^1] Power-of-two is a number of the form $2^{n}$ where n is a positive integer, negative integer, or zero. value in the range $2^{-127}$ to $2^{127}$.
Given K input elements in a source-format, $V_{i}$ (typically FP32); $1\leq i\leq K$, conversion to MX-format consists of computing X and $Q_{i}$ such that $Q_{i}\timesX$ is the decoded value corresponding to $V_{i}$. X and $Q_{i}$ are stored in memory instead of $V_{i}$. Thus, X serves as the decoding scale factor to decode quantized values back to their high-precision counterparts. This is depicted in Figure 1.
A tensor in source-format is sub-divided into blocks of K elements and converted into MX-format to be stored in memory and/or processed by math units in hardware. Tensor Cores in Blackwell consume X and $Q_{i}$ to compute the dot product of two MX-formatted blocks. If the accumulated output of a dot-product operation in Tensor Cores is FP32, then this output is thereafter quantized to MX-format if a subsequent operation consuming the output needs it in that format. The conversion process is, $Q_{i}=\verb+Quantize_to_fp8+(V_{i}/X)$. Section 3 describes the conversion details. Fine-grained scaling helps each MX block independently align to the needed range of input values before quantization.
@@
Table 1 shows the MX-formats supported in Blackwell. The data type column uses the convention ExMy to denote x-bit for floating-point exponent and y-bit for mantissa. For a fixed bit-width, floating-point numbers trade-off exponent width with mantissa width. Thus, MXFP8 E4M3 is a 8-bit data type with 1 sign bit, 4-bit for FP exponent and 3-bit of mantissa. Similarly, E5M2 is an 8-bit type with 5-bit for exponent and 2-bit for mantissa. Compared to E4M3, E5M2 can represent a larger dynamic range, but with less precision. E5M2 follows IEEE 754 conventions [6] for representation of special values, whereas the remaining data types extend the dynamic range by not representing Infinity and NaN (E4M3 has only one bit-pattern for NaN [7]). More bits in the exponent field translates to a larger range while more bits in the mantissa field translates to more precision within a given range. Every floating-point type has a dynamic range that it can represent — we denote this range in terms of binades which is the log2-ratio of the maximum to the minimum finite representable in that format.
Figure 2: Pre-training a 8B LLM on 15T tokens. Top: Training behavior of BF16 vs MXFP8. Bottom: Comparing BF16, FP8 and MXFP8’s downstream task scores on MMLU and a set of 9 reasoning tasks. MXFP8 numerics use our proposed rounding method and E4M3 for all quantized tensors.
@@ 3 Pre-training with MXFP8
In this section we review MXFP8 training results, training recipe, and the algorithm for converting to MXFP8 used for training. Ablation studies explain the recipe and conversion algorithm choices.
3.1 Training Results
To establish the equivalence of MXFP8 and BF16 for model accuracy we pretrain an 8B parameter model on 15T tokens. We observe:
  1. Validation perplexity when using MXFP8 matches that of BF16 (top plot in Figure 2). There is less than 0.50% difference between MXFP8 and BF16 validation perplexity values throughout the pre-training run.
  2. Downstream task evaluation scores match between MXFP8 and BF16 sessions (bottom plots in Figure 2).
The same equivalence is also observed for smaller models or smaller datasets, making MXFP8 a preferred option to efficiently train LLMs. The 8B parameter model [8] consists of 32 transformer blocks, 32 attention heads. Hidden size is 4096, GQA group size is 8, KV-channels count is 128, sequence length during pre-training is 8192. Initial learning rate is 6e-4 that cosine decays to 6e-6. A phased data-blending approach is used to train the model: in the first phase, a data mixture that promotes diversity in data is used and in the second phase high-quality datasets (e.g., Wikipedia) are used. We switch to the second phase at the 60% point of training. This blend style has also been used in other large scale pre-training setups [9].
Training was carried out using Megatron-LM software [10] on 3072 Hopper GPUs using batch size 768. MX operations were simulated by converting BF16 inputs to MXFP8 and back to BF16 prior to GEMM operations.
Evaluation scores are measured on two sets of downstream tasks: (1) 5-shot score on MMLU [11] and (2) Averaged 1-shot score across 9 general reasoning benchmarks: ARC-Challenge and ARC-Easy [12], Race [13], PIQA [14], Winogrande [15], Hellaswag [16], OpenBookQA [17], Social IQA [18] and Commonsense QA [19].
[…]
3.2 MXFP8 Training Recipe
The recipe is defined by two choices: which tensors to quantize to MXFP8 and which FP8 binary encoding to use for different tensors.
Figure 3: Top: Transformer layers quantized to MXFP8 inside a single transformer block. Bottom: Training workflow for a single layer during forward-pass (FPROP), error-gradient pass (DGRAD) and weight-gradient pass (WGRAD).
3.2.1 Tensors to Quantize
We execute GEMMs with weights (i.e. QKV and Proj in attention and FFN, as shown in Figure 3) in MXFP8. Therefore, activation, weight, as well as activation gradient input tensors for these GEMMs are converted to MXFP8. The Batch-Matrix Multiplications (BMM1, the query-key dot product and BMM2, the attention score-value product) in the self-attention layer, along with operations like Softmax, Act-func and residual-add are in high-precision. The input embedding layer and the final output-projection layer are also in BF16 or FP16. We leave it to future work to explore reducing the precision of these operations.
The aforementioned tensors are quantized to MXFP8 for all transformer blocks of a model, in contrast to per-tensor [9] or per-row quantization [20] for FP8, where some transformer layers were left in BF16. As a result, MXFP8 due to its fine-granularity scaling factors enables accelerating more of an LLM’s layers.
During training, with MXFP quantization, the training framework keeps two copies of each tensor, each copy is quantized along the axes of dot-product reduction (row and column). Figure 3 shows how each tensor is used in forward (FPROP), weight-gradient (WGRAD) and activation-gradient (DGRAD) computation during the training loop. Since each tensor is used in non-transposed and transposed form, quantization needs to occur along two separate axes (row and column).
@@ 3.2.1 Tensors to Quantize
3.2.2 FP8 Encoding Choice
Our MXFP8 recipe uses E4M3 encoding for all tensor types - weights, activations, and activation gradients. This is in contrast to coarser-scaled FP8 training [9, 7, 21, 22], which used E4M3 for weights and activations but E5M2 for gradients. With fine-grained scaling of MXFP8, the dynamic range requirement at a 32-element block size is sufficiently captured by 17.8 binades of E4M3 type. Once the range requirements are met, precision (or sampling) becomes important. We demonstrate this recipe choice with an empirical study on an 843M and 8B parameter models. Figure 4(a) shows that E4M3 is the preferred choice for weights and activations, as using E5M2 for these tensor types (the blue and purple curves, respectively) results in a worse perplexity. While either E5M2 or E4M3 encodings are comparably good for the 843M parameter model, in Figure 4(b) we see that the larger 8B parameter model requires the higher precision of E4M3 for activation gradients (orange curve), as E5M2 gradients (yellow curve) lead to a substantial degradation in perplexity.
3.3 Conversion from FP32 to MXFP8
Often the original high precision (for example, FP32 or BF16) values within a block are outside the representable range for the target MX format, both underflowing the minimum and overflowing the maximum representable magnitude. To address this, prior to conversion all values are multiplied with a scale factor so that no value overflows the MX format range, while minimizing flushes to 0. The shared per-block scale factor, X, is computed based on the maximum absolute value (amax) among the 32 high-precision input values, i.e. amax = $\max\|V_{i}\|;1\leq i\leq 32$. The goal is to map this amax in the input-format to be the largest representable value in the desired MX-format. Once X is computed, $V_{i}$ is divided by $X$ and the resulting value is quantized to a FP8-representable number using Round-to-nearest-ties-to-even (RN) rounding. Furthermore, saturating conversion mode is used (as described in OCP FP8 specification [23]) i.e. magnitudes exceeding FP8-maximum are clamped to respective maximum representative magnitude in quantized format. Special care is taken if some or all elements in the input are Infinity and/or not-a-number (NaN).
$\text{expX}_{\text{float}}\leftarrow\log_{2}(X_{\text{float}})$ $\triangleright$ extract the exponent of X (de-biased form)
$\text{expX}_{\text{int}}\leftarrow\text{ceil}(\text{expX}_{\text{float}})$ $\triangleright$ round-up
$X\leftarrow\text{clamp}(\text{expX}_{\text{int}},-127,127)$ $\triangleright$ clamp to min/max E8M0 representable
$X\leftarrow X+127$ $\triangleright$ add bias
return $2^{X}$ $\triangleright$ store X
Algorithm 1 outlines the scale-factor computation used by training sessions in Section 3.1. Step 3 is the key part, rounding-up the scale factor towards positive infinity, when converting X to a power-of-2 value. This is different from the scheme suggested by OCP v1 specification, which effectively rounds down the scale factor. Since during quantization to MXFP, high precision values are divided by X, rounding up X ensures that after scaling no $V/X$ value exceeds the maximum magnitude representable in the destination format. Conversely, rounding down, as in OCP v1 suggestion, will lead to scaled values overflowing the representable range, introducing more quantization noise.
We hypothesize this aspect leading to MXFP8 training issues when using OCP v1 conversion scheme as shown in Figure 5. As can be seen in the figure, these issues are eliminated when using our proposal in Algorithm1. A more detailed description of conversion workflow is provided in Appendix Sec. A.1.
Figure 5: Comparing scale factor rounding modes suggested in the OCP v1.0 specification and our work. E4M3 and E5M2 are MXFP8 formats; notation E4M3(W) denotes E4M3 data type for Weights. Similarly, E4M3(A) and E4M3 (G) refer to activations and gradients, respectively.